### v1

@@ 4.6 Ablation Experiments
In this section, we perform ablations using the DeepSeek model with a cache size of 32, evaluated on Wikitext for the perplexity/cache miss rate trade-off.
@@
As discussed in Section 4.1, we always select Top-J experts ($J<K$) even if they are not in cache. Figure 4 shows the effect of varying $J$ on model performance. While baseline Max-rank and Cumsum-threshold methods clearly benefit from keeping Top-J intact, impact on our Prior method is less pronounced. It leads to improved perplexity for certain models (QWEN, Phi) while being detrimental to other models (DeepSeek, Mixtral). Since keeping $J>0$ is critical to baseline methods while providing good-enough trade-off between perplexity and cache miss rate for our Prior method, we decide to keep it fixed per-model across all methods.
It is important, however, to note that different values of $J$ optimize different constraints. For example, $J=0$ enables very low ($<2\%$) cache miss rates while, potentially, damaging language model performance. In practice, flexibility of our training-free approach allows $J$ to be adapted easily for deployment on different targets. Therefore, one can choose $J$ depending on the specifics of the target setting.
@@
Figure 8: Performance on WikiText for different logits estimation strategies
5 Related Work
Speculative Routing and Weight Preloading: Recently, various approaches have been developed to optimize MoE inference through speculative routing and weight preloading Cui et al. (2023); Xue et al. (2024); Eliseev & Mazur (2023); Du et al. (2024). These methods primarily focus on expert prefetching strategies to boost cache hit rates. For example, MoE-Infinity Xue et al. (2024) uses look-ahead routing and caches the most frequently used experts. SiDA-MoE employs an offline-trained hash function to reduce expert selection overhead. However, SiDA-MoE’s approximate caching results in a significant drop in perplexity, underscoring the difficulty of maintaining accuracy. Additionally, Eliseev & Mazur (2023) combines lookahead routing and LRU caching, necessitating the pre-loading of multiple experts well in advance to optimize the overlap between data movement and inference calculations. While this approach helps conceal memory transfer costs, it can adversely affect model accuracy. Speculative routing to improve cache hit-rate without loss of accuracy is particularly challenging on memory-constrained devices where token generation is more memory-bound than compute-bound. In contrast, our method enhances routing consistency and exploits locality priors, achieving minimal changes in perplexity and significant latency improvements. Our approach is orthogonal to speculative routing methods.
Efficient LLM Inference: Techniques such as LLM quantization Dettmers et al. (2022); Frantar et al. (2023); Shao et al. (2024) and compression Frantar & Alistarh (2023) can significantly reduce the memory footprint of large language models (LLMs). Recent studies have explored pruning or merging MoE experts based on their utilization to conserve memory Chen et al. (2022); Li et al. (2024); Lu et al. (2024), though this can result in decreased performance. To facilitate the deployment of off-the-shelf MoEs on devices with limited memory without a model surgery that would change the architecture, previous works have focused on optimizations like offloading parameters to host memory Alizadeh et al. (2024); Yu et al. (2024); Jung et al. (2023); Rasley et al. (2020). These approaches employ dynamic scheduling strategies to offload sparse parameters while maximizing the overlap between expert loading and computation. For instance, DeepUM Jung et al. (2023) records CUDA kernel execution patterns to predict which kernel is likely to be executed next. However, MoE model execution patterns can vary significantly depending on the task and context. Our proposed method enhances expert cache hit rates without requiring complex hardware implementations and remains agnostic to the specific task being addressed.
6 Conclusion
In this work, we introduced a novel training-free cache-aware expert routing approach for Mixture of Experts (MoE) models, tailored to enhance inference efficiency on memory-constrained hardware. By increasing cache hit rates through consistent routing, our method achieves a significant reduction in latency while preserving model accuracy, demonstrating up to a 2$\times$ speedup over regular routing with LRU caching techniques on mobile devices.
Our method is highly adaptable, requiring no model retraining, making it suitable for a wide range of hardware configurations. This adaptability reduces deployment costs and facilitates the practical application of MoE models in real-world, resource-limited environments.
Appendix
Appendix A LRU Throughput
In Figure 9 we show on-device performance for various cache sizes. Similar to Section 4.4, we use best LRU performance as a reference point of 1×1$\times$1 ×.
One can clearly see that increasing cache size beyond $30$ in case of 4-bit quantized model on 12GB device (left) and beyond $45$ in case of 8-bit quantized model on 16GB device (right) leads to decrease in performance. This is due to the fact that increasing cache size reduces available memory, causing the OS to reload uncached components (e.g., kv-cache, activations) for each token, which significantly slows down inference.
Given these results, we use same cache size of $30$ for 4B and $45$ for 8B in experiments in Section 4.4.
Appendix B Smaller cache sizes
Throughout all our experiments in the paper, we use a cache size that is $\frac{1}{2}$ times the total number of experts. In Figure 12 we demonstrate performance of our approach for models with cache size equal to $\frac{1}{4}$ of the total amount. Our Prior method is outperforming all other methods across the board. Note that it is doing so without any changes to method itself making it appealing for final user who might want to deploy it across a range of devices.
Appendix C LRU Ablation
Figure 10: Wikitext Perplexity results for varying LRU addition approaches.
Figure 11: Prior vs Learned Prior performance.
In Figure 10 we show that our approach (Top-Bottom) for applying the LRU cache, as described in Section 4.6 performs better than Bottom-Top approach. However, the difference is fairly insignificant and thus the LRU strategy should not be considered an important factor.
Appendix D Learned Prior
@@
However, it is also possible to make the proposed method learnable. To do so, we propose learning the additive prior term using a small cache MLP layer conditioned on the cache state. Specifically, we implement this by employing a two-layer MLP that takes both the cache state and router logits as input, outputting a single bias vector. This bias is then added to the router logits and used for the next step. The loss function is optimized on the softmax outputs. We force weights assigned to experts that are in the cache but not in Top-K to move closer to one while penalizing weights for those that are in Top-K but not in the cache to move closer to zero.
A downside of learnable priors is that we are required to train many independent cache MLPs to obtain trade-off points between cache hit rate and model accuracy. We conducted an experiment to verify the effectiveness of learning this bias but did not observe notable improvements. This optimization is challenging and the results were close to the prior performance without outperforming it. Moreover, this approach would be less beneficial in settings with multiple deployment targets. Therefore, the training-free nature of our approach proves very advantageous for real-world applications.

### v2

@@ 4.8 Ablation Experiments
In this section, we perform ablation experiments on the WikiText dataset to demonstrate the general applicability of our method across a range of cache sizes, compare its performance to the theoretical bounds of the optimal cache eviction policy, and justify our hyperparameter choices.
Optimal Cache Eviction
As a theoretical upper bound for cache policy, we consider Belady’s algorithm (Belady, 1966), which evicts the expert that will be needed farthest in the future. This lossless policy requires perfect foresight of future router decisions, making it unattainable in practice but useful as an upper bound on cache hit rate. Figure 10 shows that while Belady’s algorithm roughly halves cache miss rates compared to LRU.
Conventional caching methods (e.g., LRU, Belady’s optimal) are lossless and thus cannot surpass the theoretical cache hit rate bound set by the model’s own routing. In contrast, our approach introduces controlled, lossy, reranking of experts, allowing the model to swap lower-ranked experts for in-cache ones, at the cost of a small, tunable increase in perplexity. This design enables us to trade a negligible amount of accuracy for a substantial gain in cache efficiency, even surpassing the oracle bound. As can be seen in Figure 10, our Cache-Prior method achieves similar or even lower miss rates with only minimal perplexity increases.
Cache Size
We investigate whether our findings hold across different cache sizes, beyond our default of half the total expert count. Figure 11 shows results for cache sizes ranging from one expert to all experts, comparing LRU, optimal caching, and our Cache-Prior method (with an LRU cache). Unlike the first two approaches, Cache-Prior introduces a perplexity trade-off controlled by $\lambda$. To standardize this comparison, we select the $\lambda$ value for each model and cache size that minimizes cache miss rate while keeping the perplexity increase within specific thresholds (1%, 5%, and 10%, shown in different shades of purple).
As expected, when the cache size equals $N$, all experts fit in memory, and the miss rate drops to zero for all methods. For the optimal cache baseline, reducing cache size to half the number of experts improves performance over LRU, but the benefit diminishes as cache size shrinks further, becoming negligible for cache sizes $\leq k$.
In contrast, our Cache-Prior method consistently enhances performance, particularly at smaller cache capacities. At the extreme of a cache size of one, even the optimal caching strategy yields only marginal improvements. For models with high values of $k$, our method behaves similarly, as cache reuse still results in $k-1$ cache misses. However, for models with smaller values of $k=2$, such as Mixtral and Phi, Cache-Prior significantly improves cache hit rates, achieving up to a 20% reduction in cache misses even at the limit of cache size = 1.
Remarkably, with just a 1% increase in perplexity, our Cache-Prior method outperforms the theoretical optimal caching bound for all cache sizes and models, except Phi-3.5-MoE-Instruct. When allowing a 5% increase in perplexity, our method surpasses the optimal bound across all models and cache sizes. This demonstrates the effectiveness of Cache-Prior, as it can exceed the theoretical optimal caching performance with only a minimal trade-off in downstream task performance.
Figure 11: Cache size ablation on Wikitext. Cache size of half of the total number of experts $N$ (highlighted in bold) is used throughout the rest of this paper. Cache size of $k$ is highlighted in blue.
Impact of Initial Cache State
A potential concern with cache-aware expert routing is that the initial cache content might bias the model to repeatedly select the same experts, thereby limiting diversity and adaptability in expert selection. To investigate this, Figure 19 in Appendix G visualizes cache hits/misses for a GSM8K sample at the start of prompt encoding, for the original routing and our Cache-Prior ($\lambda=0.5$ and $\lambda=0.8$) method. We consider two scenarios: (1) the initial cache is empty and fills as experts are selected during processing of the first few tokens, and (2) the cache is initialized with a random set of experts. We provide more details on this ablation in Appendix G.
Our results show that, for moderate values of $\lambda$ (e.g., 0.5), the initial cache state has minimal long-term impact: after processing a few tokens, the distribution of expert activations and the cache state converge, regardless of initialization. This indicates that the routing mechanism does not remain biased toward the initial cache content, and the method remains robust to different initial cache states without degrading model diversity or generation performance. In practice, the strength of the bias towards cached experts can be controlled by the parameter $\lambda$. When using excessively larger $\lambda$ values (e.g., 0.8), the model may become overly reliant on the cached experts, potentially reducing predictive performance.
@@
As discussed in Section 4.2, we first select the top-$J$ experts ($J<K$) from the router’s normal top-$K$ selection, then complement the remaining experts by favoring those already in the cache, ensuring both critical experts and cache-efficient choices are included. Figure 4 shows the effect of varying $J$ on language modeling performance. While baseline methods like Max-rank and Cumsum-threshold experience a significant performance drop without loading the top-$J$ experts, Cache-Prior is more robust to changes in $J$. The impact on Cache-Prior is model-dependent, with minor improvements in perplexity for models like DeepSeek, Mixtral, and Phi, and a slight decrease for models like QWEN. Given that $J>0$ benefits baseline methods and produces adequate results for Cache-Prior, we keep it fixed across all methods.
5 Related Work
Expert Caching
Caching is a common optimization in MoE models, given their high memory requirements. Most prior works rely on a cache with an LRU eviction policy (Eliseev & Mazur, 2023; Yuan et al., 2024b; Yi et al., 2023; Hwang et al., 2024; Huang et al., 2023), while others adopt a least-frequently-used (LFU) eviction policy (Xue et al., 2024a) or evict experts based on more dynamic and custom measures of expert importance (Kamahori et al., 2024; Kong et al., 2023).
These methods manage expert storage and eviction without changing the model’s routing decisions or outputs and thus also maintain identical performance. However, this constraint limits their effectiveness, as they cannot change which experts are selected. To establish a performance upper bound, we include an optimal cache baseline that assumes perfect foresight of future router decisions. While this is unattainable in practice, we show that our method can surpass this theoretical bound with only a minimal increase in perplexity. Unlike conventional caching strategies, our approach introduces a controlled trade-off between accuracy and latency, enabling more flexible resource management.
Other lossy caching strategies include Adapt-MoE (Zhong et al., 2024), which dynamically adjusts cache sizes per layer, and HOBBIT (Tang et al., 2024), which decides whether to fully load, skip, or use lower-bitwidth versions of experts based on router confidence. However, both rely on complex prefetching mechanisms. In contrast, our method is much simpler, requiring only a single hyperparameter to balance accuracy and latency effectively.
Speculative Routing and Weight Prefetching
Several approaches optimize MoE inference by prefetching expert weights and using speculative routing (Cui et al., 2023; Xue et al., 2024b; Eliseev & Mazur, 2023; Du et al., 2024). These methods improve cache hit rates by predicting which experts will be needed in advance. For example, MoE-Infinity (Xue et al., 2024b) uses look-ahead routing and caches the most frequently used experts. SiDA-MoE employs an offline-trained hash function to reduce expert selection overhead, though at a significant cost to perplexity. Similarly, Eliseev & Mazur (2023) combines look-ahead routing with LRU caching, requiring the pre-loading of multiple experts well in advance to optimize the overlap between data movement and inference calculations. While this approach helps conceal memory transfer costs, it can adversely affect model accuracy. Speculative routing to improve cache hit-rate without loss of accuracy is particularly challenging on memory-constrained devices where token generation is more memory-bound than compute-bound. In contrast, our method enhances routing consistency by exploiting locality priors, achieving minimal increase in perplexity and significant latency improvements. Our approach is orthogonal to speculative routing methods.
Efficient LLM Inference
Techniques such as LLM quantization (Dettmers et al., 2022; Frantar et al., 2023; Shao et al., 2024) and compression (Frantar & Alistarh, 2023) can significantly reduce the memory footprint of large language models (LLMs). Recent studies have explored pruning or merging MoE experts based on their utilization to conserve memory (Chen et al., 2022; Li et al., 2024; Lu et al., 2024), though this can result in decreased performance. To facilitate the deployment of off-the-shelf MoEs on devices with limited memory without a model surgery that would change the architecture, previous works have focused on optimizations like offloading parameters to host memory (Alizadeh et al., 2024; Yu et al., 2024; Jung et al., 2023; Rasley et al., 2020). These approaches employ dynamic scheduling strategies to offload sparse parameters while maximizing the overlap between expert loading and computation. For instance, DeepUM (Jung et al., 2023) records CUDA kernel execution patterns to predict which kernel is likely to be executed next. However, MoE model execution patterns can vary significantly depending on the task and context. Our proposed method enhances expert cache hit rates without requiring complex hardware implementations and remains agnostic to the specific task being addressed.
6 Conclusion
In this work, we introduced a novel training-free, cache-aware expert routing approach for Mixture of Experts models, designed to improve inference efficiency on memory-constrained hardware. By increasing cache hit rates through consistent routing, our method significantly reduces latency while preserving model accuracy, achieving up to a 2×2$\times$2 × speedup over standard LRU caching on mobile devices.
Our experiments demonstrate that models remain robust to expert swapping, particularly when the top-1 expert is preserved. This suggests new possibilities for balancing performance with cache efficiency. To explore this, we introduced three routing strategies, with our Cache-Prior method proving particularly effective. Unlike prior work, which is fundamentally constrained by the optimal oracle cache bound, we show that allowing minimal deviations in expert selection enables us to surpass this theoretical limit while maintaining strong downstream task performance.
We conducted a comprehensive analysis across four state-of-the-art models—DeepSeek-V2-Lite, Qwen1.5-MoE-A2.7B, Phi-3.5-MoE, and Mixtral-8$\times$7B—on three tasks: WikiText, MMLU, and GSM8K. Our Cache-Prior method achieves a substantial reduction in cache miss rates across all models, decreasing them by more than 50%, while incurring only a minimal increase in perplexity of 0.1%–3% on the language modeling task and a negligible accuracy drop of less than 0.1% on downstream tasks. Additionally, our findings indicate that granular MoE models, such as DeepSeek-V2-Lite and Qwen1.5-MoE-A2.7B, are inherently more cache-efficient. This trend suggests a promising future for newer MoE architectures like DeepSeek v3 and Qwen2.5max, which, despite being too large for mobile deployment today, could benefit from their cache-friendly structure.
Our method is highly adaptable, requiring no model retraining, making it suitable for a wide range of hardware configurations. This adaptability reduces deployment costs and facilitates the practical application of MoE models in real-world, resource-limited environments. A key feature of our approach is that it enables explicit control over the trade-off between cache efficiency and predictive performance via the $\lambda$ parameter. While excessively large values of $\lambda$ can introduce a brief, transient bias toward the cache state and may encourage overuse of cached experts, this trade-off is tunable and can be adjusted to suit specific deployment needs. In practice, moderate values of $\lambda$ provide substantial cache efficiency gains with minimal impact on accuracy.
In summary, our cache-aware expert routing approach not only enhances inference efficiency but also opens new avenues for deploying large MoE models in resource-limited environments. This work paves the way for more practical and cost-effective applications of MoE models, ensuring their scalability and adaptability in diverse hardware configurations.
Appendix
Appendix A Optimal Expert Selection
To investigate the accuracy of the router’s expert predictions, we conduct an experiment on Wikitext using Mixtral-8x7B-v0.1. Iterating over layers, we fix the top-1 expert and systematically test all possible choices for the second expert, finding the one that minimizes perplexity. The results, shown in Figure 12, reveal that the router selects the optimal top-2 expert only 28% of the time on average. Although deeper layers improve this accuracy, the best-performing router only correctly predicts the top-2 expert in just 38% of cases. This suggests that while selecting the highest-ranked experts is important, router predictions are often suboptimal, leaving room for flexibility in expert selection without significantly impacting performance.
Figure 12: Agreement between router predictions and optimal expert selection across different layers of Mixtral-8x7B-v0.1. The y-axis represents the frequency with which a predicted expert was actually the optimal choice based on Wikitext perplexity.
Appendix B Intuitive Explanation of the Max Rank and Cumulative Sum Threshold Baselines
To provide further intuition, the Max Rank approach can be understood as follows. Rather than always selecting the top-$K$ experts with the highest router probabilities, we consider a slightly larger pool of candidate experts, specifically, those ranked in the top $M$ by the router. Among these, we give priority to experts that are already present in the cache, promoting them to the front of the ranking. This increases the likelihood that the selected experts are already loaded in fast memory, reducing cache misses and improving inference efficiency. However, to avoid degrading model performance, we do not promote cached experts that fall outside the top $M$ router ranks, ensuring that only reasonably probable experts are considered. In effect, this method balances the trade-off between cache efficiency and predictive accuracy by flexibly reordering the selection within a controlled rank window.
Figure 13: Illustration of the Max Rank and Cumulative Sum Threshold Baselines.
To provide an example, suppose the router produces a ranked list of experts $\mathbf{r}=[E_{1},E_{2},E_{3},E_{4},E_{5},E_{6}]$, and the cache currently contains $C=\{E_{3},E_{4},E_{6}\}$ (See Figure 13a for an illustration). Let $M=4$, $K=2$, and $J=1$. The top $M$ experts are $[E_{1},E_{2},E_{3},E_{4}]$. The intersection with the cache yields $[E_{3},E_{4}]$. Applying the promotion operation, we reorder the list to obtain $[E_{3},E_{4},E_{1},E_{2},E_{5},E_{6}]$. To ensure the top-J is always included, we move them to the front in order while preserving the rest $[E_{1},E_{3},E_{4},E_{2},E_{5},E_{6}]$. The top-$K$ experts selected for this token would then be $E_{1}$ and $E_{3}$.
The procedure for expert swapping in the Cumulative Sum Threshold approach is based on the same principle, however, the maximum rank $M$ is dynamically determined by thresholding the cumulative sum of sorted router probabilities. As can be seen in Figure 13b, for flat distributions, $M$ is larger and for peaky distributions, $M$ takes smaller values.
Appendix C On-Device LRU Throughput
In Figure 14 we show on-device performance for various cache sizes. Similar to Section 4.5, we use best LRU performance as a reference point of 1×1$\times$1 ×. As can be seen, increasing cache size beyond $30$ in case of 4-bit quantized model on 12GB device (left) and beyond $45$ in case of 8-bit quantized model on 16GB device (right) leads to decrease in performance. This is due to the fact that increasing cache size reduces available memory, causing the OS to offload uncached components (e.g., KV-cache, activations) for each token, requiring reloading them from flash which significantly slows down inference.
Given these results, we use same cache size of $30$ for 4-bit and $45$ for 8-bit quantized models in experiments in Section 4.5.
Appendix D Smaller Cache Sizes
Throughout all the experiments in the paper, we use a cache size that is $\frac{1}{2}$ times the total number of experts. In Figure 15 we demonstrate performance of our approach for models with cache size equal to $\frac{1}{4}$ of the total amount. Our Cache-Prior method is outperforming all other methods across the board. Note that it is doing so without any changes to method itself making it appealing for final user who might want to deploy it across a range of devices. An aggregated overview of additional cache sizes can be found in Figure 11 in the main text.
@@ Appendix D Smaller Cache Sizes
Figure 16: Effect of different strategies for estimating the logit range.
Figure 17: Performance of our learned Cache-Prior strategy on Wikitext.
Appendix E Learned Cache-Prior
@@
However, it is also possible to make the proposed method learnable. To do so, we propose learning the additive cache-prior term using a small cache MLP layer conditioned on the cache state. Specifically, we implement this by employing a two-layer MLP that takes both the cache state and router logits as input, outputting a single bias vector. This bias is then added to the router logits and used for the next step. The loss function is optimized on the softmax outputs. We force weights assigned to experts that are in the cache but not in top-$K$ to move closer to one while penalizing weights for those that are in top-$K$ but not in the cache to move closer to zero.
A downside of learnable cache-priors is that we are required to train many independent cache MLPs to obtain trade-off points between cache hit rate and model accuracy. We conducted an experiment to verify the effectiveness of learning this bias but did not observe notable improvements. This optimization is challenging and the results were close to the cache-prior performance without outperforming it. Moreover, this approach would be less beneficial in settings with multiple deployment targets. Therefore, the training-free nature of our approach proves very advantageous for real-world applications.
@@
Appendix F Impact of Prompt Length on Throughput
Figure 18 shows the influence of prompt length (short: 40-60 tokens; long: 300-400 tokens) on throughput for the Qwen1.5-MoE-A2.7 model deployed with a cache size of 30 experts. Across nearly all values of $\lambda$, longer input prompts yield higher throughput.
Figure 18: Influence of prompt length on relative throughput across varying $\lambda$ values, for the Qwen1.5-MoE-A2.7 running on the 16GB device with a cache size of 30 experts.
Appendix G Impact of Initial Cache State
In this section, we qualitatively assess how the initial cache state affects expert selection. Figure 19 visualizes cache hits and misses for a GSM8K sample during the prompt encoding phase for the Phi-3.5-MoE-instruct model. We compare both the original routing and Cache-Prior methods for a cache size of 8 experts across four scenarios: (1) original routing with an empty cache, (2) Cache-Prior ($\lambda=0.5$) with an empty cache, (3) Cache-Prior ($\lambda=0.5$) with a randomly initialized cache, and (4) Cache-Prior ($\lambda=0.8$) with the same random cache state as in (3). This setup allows us to examine the influence of both the initial cache content and the $\lambda$ parameter on expert selection. In all these setting, we set $J=1$, for guaranteed top-1 expert loading.
Our results show that, for moderately large values of $\lambda$ (e.g., 0.5), the initial cache state has minimal long-term effect: after processing a few tokens, the distribution of expert activations and the cache state converge, regardless of how the cache was initialized. This suggests that the routing mechanism is not persistently biased toward the initial cache content. However, with excessively larger $\lambda$ (e.g., 0.8), we observe that the model tends to overly reuse experts in the cache, which may negatively impact predictive performance.
Appendix H Qualitative On-Device Results
Table 2 presents sample text from our on-device autoregressive text generation experiment. It is evident that incorporating our Cache-Prior method has minimal impact on the quality of the generated results. For more details, refer to Section 4.5 of the main text.